OpenStudy (anonymous):

Which of the point-slope equations below is correct for the line that passes through points (-1,-4) and (3,2)? Check all that apply.

OpenStudy (anonymous):

A. y-1=3/2(x+3) B. y+4=3/2(x-1) C. y+4=3/2(x+1) D. y-2=3/2(x-3) E. y+4=2/3(x+1)

OpenStudy (anonymous):

@anonymous97 what is this

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!}\) Plug these points into the formula y2-y1/x2-x1